OpenStudy (anonymous):

which of the following is equivalent to the expression : (cotθ-1)/(1-tanθ) a. cotθ b. cscθ c. secθ d. 1 e. tanθ

OpenStudy (anonymous):

Re-write cot theta as cos/sin and tan as sin/cos to start. Then take the ones and get a common denominator.

OpenStudy (anonymous):

\[\frac{ \cot \theta - 1 }{ 1 - \tan \theta } = \frac{ \frac{ \cancel{\cos \theta - \sin \theta} }{ \sin \theta } }{ \frac{ \cancel{\cos \theta - \sin \theta} }{ \cos \theta } } =\frac{ \cos \theta }{ \sin \theta }\]

OpenStudy (anonymous):

All good now, @lamp12345 ?

OpenStudy (anonymous):

i think, thanks

OpenStudy (anonymous):

Ask any question you want if you're not sure.

OpenStudy (anonymous):

I got cotθ. is that right?

OpenStudy (anonymous):

Yes that is right, but you should know your identities a bit better.

OpenStudy (anonymous):

When you see tan, you should automatically see sin/cos in your head. Similarly with cot = cos/sin. If you don't, you'll have to study those more or you'll never be able to do this type of problem.

OpenStudy (anonymous):

alright thanks!

OpenStudy (anonymous):

np! Have a great day! Nice working with you! @lamp12345